OpenStudy (anonymous):

Algebraically show that n P r = (n C r) X (r P r) and describe the significance of this relationship in your own words.

OpenStudy (anonymous):

nPr = n! / ( n-r ) ! nCr = n! / [ (n-r)! r! ] r P r = r! / ( r-r)! , and 0! = 1, n! = n* ( n-1)!

OpenStudy (anonymous):

so putting it together we have nPr = n! / ( n-r ) ! nCr x rPr = n! / [ (n-r)! r! ] *r! / ( r-r)!